OpenStudy (anonymous):

What is the exact volume of the cylinder? A. 224(pi) mm³ B. 3136(pi)mm³ C. 3584(pi)mm³ D. 6272(pi)mm³

OpenStudy (anonymous):

Radius-14 Height-16

OpenStudy (k_lynn):

The volume formula of a cylinder is \[V=\Pi r ^{2}h\]R being the radius and h being the height. So all you have to do is plug the numbers in for the variables and solve the equation.

OpenStudy (k_lynn):

Pi is also equivalent to 3.14.

OpenStudy (k_lynn):

So, once the numbers are correctly substituted, it should look like this \[3.14\times14^{2}\times16\]OR....\[3.14\times14\times14\times16\]

OpenStudy (k_lynn):

So what do you get when you do that math @doresmith

OpenStudy (anonymous):

I got 9847.04 but its not one of the answer choices.

OpenStudy (k_lynn):

I got the same thing. What are the answer choices?

OpenStudy (anonymous):

Is it B? cause what i did is the answer choices it also has the pi sign which probably have to multiple that with 3.14 which is the same answer as the one that we got.

OpenStudy (anonymous):

3136*3.14=9847.04

OpenStudy (k_lynn):

If it has the pi sign in the options, then all you do is not multiply pi in the equation. So... \[14^{2}\times16\]

OpenStudy (k_lynn):

Then once you get that answer, you put the pi sign back in \(\pi\)

OpenStudy (anonymous):

so which would be the answer choice?

OpenStudy (k_lynn):

All you have to do now, is finish that math: \[14\times14\times16\] and then put \(\pi\) in front of it. Then you have your answer.

OpenStudy (anonymous):

So the correct answer is B?

OpenStudy (k_lynn):

What were your options?

OpenStudy (anonymous):

A. 224(pi) mm³ B. 3136(pi)mm³ C. 3584(pi)mm³ D. 6272(pi)mm³

OpenStudy (k_lynn):

Yep, it was B. :D Great job!

OpenStudy (anonymous):

Thank you for helping. I gave you a medal and became a fan.